OpenStudy (anonymous):

for what value is this rational expression undefined? x+2/2x-1

OpenStudy (anonymous):

0

OpenStudy (anonymous):

please use brackets

OpenStudy (anonymous):

Hi @cyter - please don't provide 'just an answer' - we're here to explain/teach, thanks! Oh, and welcome back @rememberwhen2012 - long time no see ;)

OpenStudy (hoblos):

a rational expression is undefined when the denominator =0 2x-1=0 x=1/2

OpenStudy (anonymous):

is it (x+2)/(2x-1) or (x+2)/x -1

OpenStudy (anonymous):

if the denominator is 0, it is undefined x=1/2

OpenStudy (anonymous):

thankyou everyone!

OpenStudy (anonymous):

There is a difference between \[ \huge x+2/2x-1\] and \[\huge \frac{x+2}{2x-1}\] In the first case 2/2x would be undefined for x=0. 2*0=0; but 2/0 is undefined. It doesn't make sense so every fraction with zero in the denominator is undefined and invalid. in the second case x=1/2 would lead to it being undefined.

OpenStudy (anonymous):

And then @cyter TOTALLY redeems himself (said in Dumb&Dumber voice)! Great work cyter, and thank you, really appreciate your work!

OpenStudy (anonymous):

@cshalvey Thank you for the great work you are doing. You are the people who make openstudy unique. Thanks.